OpenStudy (anonymous):

how to solve x^ - 2x + 1 = 0 in quadratic formula

OpenStudy (kidrah69):

What are your a b c values?

OpenStudy (anonymous):

well the a = 1 the b is -2 and c is 1 @kidrah69

OpenStudy (kidrah69):

\[x=\frac{ -(-2) ±\sqrt{(-2)^2-4(1)(-2)}}{ 2(1)} \] solve.

OpenStudy (kidrah69):

Solve what is in the square root sign but DO NOT square root anything!

OpenStudy (anonymous):

ok . im doing it rn @kidrah69

OpenStudy (kidrah69):

:) let me know one you are done buddy :)

OpenStudy (anonymous):

ok i got x = -5 and x = 7 @kidrah69

OpenStudy (kidrah69):

as final answers?

OpenStudy (anonymous):

oh wait no i have to branch it out right like x - 5 = 0 ?

OpenStudy (kidrah69):

oh wait oops >_<

OpenStudy (kidrah69):

im sorry! i put -2 in b's place

OpenStudy (kidrah69):

*a's place

OpenStudy (kidrah69):

\[x=\frac{ -(-2) ±\sqrt{(-2)^2-4(1)(1)}}{ 2(1)} \]

OpenStudy (anonymous):

no u put it in c haha ok let me resolve then @kidrah69

OpenStudy (anonymous):

so (x -1) and (x - 1)

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

yes! :D

OpenStudy (kidrah69):

now set it equal to 0

OpenStudy (kidrah69):

x-1=0

OpenStudy (anonymous):

no i already did tht haha @kidrah69

OpenStudy (kidrah69):

whats the final answer :) ?

OpenStudy (anonymous):

the one i just told you was haha @kidrah69 because tht was after i already put it to the 0

OpenStudy (kidrah69):

No you still need to set it equal to 0 (x-1)^2 just equals 1 so x=1

OpenStudy (anonymous):

no ok. so like i got (x + 1) (x +1) then i branch them out and solve : x + 1 = 0 and x + 1 = 0 and my final answer is (x-1)(x-1)

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

no :/

OpenStudy (kidrah69):

(x -1) and (x - 1) x-1=0 +1 +1 x=1

OpenStudy (anonymous):

what.. show me !! haha @kidrah69

OpenStudy (kidrah69):

http://www.mathwarehouse.com/quadratic/quadratic-formula-calculator.php look at this :P its proof!!! haha

OpenStudy (anonymous):

show me how you did the whole thing . i dont want to do it off a calculator lol @kidrah69

OpenStudy (kidrah69):

did u even look at the link? -_-

OpenStudy (anonymous):

yeah i did. and its a calculator haha @kidrah69 dnt be lazy >.<

OpenStudy (kidrah69):

http://prntscr.com/78z689

OpenStudy (kidrah69):

well the work is done :P

OpenStudy (anonymous):

haha whatever :P @kidrah69

OpenStudy (kidrah69):

ey i know how to do this :P its just i was making sure my work was right and i realized i made a mistake :P

OpenStudy (anonymous):

still smarter than me c: @kidrah69

OpenStudy (kidrah69):

nawh :3